# FG-PE: Factor-graph Approach for Multi-robot Pursuit-Evasion

### v1

@@ Abstract
Index Terms:
multi-robot system, factor graph, pursuit evasion, uncertainty, tracking
I Introduction
@@
Several methods have been utilized to solve the PE problem. Classical methods use graph-based techniques, treating the PE problem as a search problem [6]. Reinforcement learning (RL) approaches [7, 8] and game theory methods [9] have also been investigated. However, existing methods often lack an efficient representation of the relationships between pursuers, evaders, and their environment. For instance, if one pursuer loses its history of movements, retrieving that information can be challenging. Utilizing a graph-based approach can establish structural connections between entities, improving the overall understanding of the situation. In addition, the aforementioned methods do not account for uncertainties in predictions and do not facilitate message passing among robots. This limitation affects the scalability and versatility of PE variants.
@@
In this paper, we present FG-PE, a method that employs factor graphs (FG) [10] to tackle the PE problem. Factor graphs provide a flexible framework capable of adapting to varying numbers of pursuers and obstacles. This approach addresses prediction uncertainty by guiding pursuers to minimize capture time, distance, and uncertainty, utilizing information from sensor measurements and past observations. The main contributions of the proposed method are: (1) FG-PE provides a factor graph-based solution for addressing the pursuit-evasion problem. By constructing a factor graph, the proposed method can estimate the position of the evader at each time step. The generated graph enables querying the position of each pursuer and other entities present in the scene. To the best of our knowledge, this paper is the first to formulate the PE problem as a factor graph. (2) FG-PE can estimate the evader’s position and plan the pursuers while considering short time windows. This means that at each time step, it can determine the planned position for each pursuer for the subsequent time step. (3) FG-PE accounts for uncertainty in predictions using the probabilistic nature of factor graphs. It aims to minimize this uncertainty by providing a plan through factor graphs, which enable efficient representation and optimization in the presence of noise. Videos of the experiments are available on the website of the project: https://sites.google.com/view/pursuit-evasion. (4) FG-PE is easily extendable to multiple pursuers and allows for high obstacle-wise scalability. (5) FG-PE is robust against lost messages and performs well even when some messages are dropped within the graph structure.
The rest of the paper is organized as follows: Sec. II presents background. Sec. III introduces the proposed method. Sec. IV shows the experimental results. Sec. V concludes the paper.
II Literature Review and Background
@@
The PE problem finds applications in various robotics scenarios, such as surveillance, search and rescue operations, and multiplayer games [11]. PE involves two groups of agents: pursuers and evaders. The evaders aim to reach a predefined goal, such as a specific location, object, or target area, while the pursuers try to prevent them from doing so. Key aspects of the PE problem include multi-agent tracking [12] and path planning [2]. There are three main approaches solving the PE problem: search-based, learning-based, and graph-based approaches.
@@
A direct method for addressing the PE problem is to use search algorithms like Depth-First Search (DFS) [13] or Breadth-First Search (BFS) [14]. A critical aspect of the pursuit-evasion problem is the speed of the agents. Furhan et al. [15] introduced a method that accounts for uncertainties in the pursuers’ speed, which can lead to issues such as moving to unintended locations.
@@
There are several real-world applications that require the collaboration of multiple robots such as object transportation [23] and task allocation [24]. One approach to coordinating the necessary communication among robots is by utilizing graph representation. Factor graphs [10], described in Sec. III-C in detail, can represent complex relationships between agents concisely and are suitable for addressing a wide range of problems such as kinodynamic motion planning that fully considers whole-body dynamics and contacts [25]. Factor graphs have been widely used to solve various multi-robot problems, such as localization [26], path planning [27], and control [4]. In this paper, we utilize factor graphs to address the PE problem. To the best of the authors’ knowledge, there is no existing work, that formulates the PE problem using a factor graph representation.
III Proposed Method
In this section, the problem statement and the proposed solution are presented. Table I lists the notations used.
Var. Description
@@ III Proposed Method
$f^{mp_{j}}_{t}$ Measurement factor btw $j^{th}$ pursuer and the evader at time $t$
$f^{m_{j}o_{i}}_{t}$ Measurement factor btw $j^{th}$ pursuer and $i^{th}$ obstacle at time $t$
$f^{p_{j}mov}_{t}$ Planning factor of the $j^{th}$ pursuer at time $t$
$f^{cp_{j}}_{t}$ Collision avoidance factor for $j^{th}$ pursuer at time $t$
$f^{o_{i}p_{j}}_{t}$ Obstacle avoidance factor btw $j^{th}$ pursuer and $i^{th}$ obstacle at time $t$
$e^{qd}_{t}$ Cost of dynamic factor of the evader
$e^{p_{j}d}_{t}$ Cost of dynamic factor of the $j^{th}$ pursuer
$e^{mp_{j}}_{t}$ Cost of measurement factor btw $j^{th}$ Pursuer and Evader
$e^{m_{j}o_{i}}_{t}$ Cost of measurement factor btw $j^{th}$ pursuer and $i^{th}$ obstacle
$e^{cp_{j}}_{t}$ Cost of collision avoidance btw pursuer $j$ and other pursuers
$e^{o_{i}p_{j}}_{t}$ Cost of obstacle avoidance factor btw pursuer $j$ and obstacle $i$
TABLE I: Variables used in this work.
III-A Problem Statement
@@
$\mathcal{U}=\{v\in\mathbb{R}^{2}\mid||v||\leq v^{\text{max}}\}$ (1)
=$\mathcal{R}=\{\omega\in\mathbb{R}^{2}\mid||\omega||\leq\omega^{\text{max}}\}$ (2)
@@
III-B Path Planning of Evader
The evader aims to reach a predefined goal, known only to itself. The evader employs the Dynamic Window Approach (DWA), a popular method in robotics for motion planning and obstacle avoidance [28]. DWA is effective in scenarios where a robot navigates through moving obstacles. It dynamically adjusts the robot’s velocity and steering commands based on real-time sensory input and desired goal states. By continuously evaluating and updating potential future trajectories within a dynamically defined window, the evader can efficiently navigate toward its goal while avoiding obstacles and pursuers. The evader iterates through all possible trajectories based on angular velocity and linear velocity, selecting the best movement for the next time step. In this paper, we utilize a modification of the DWA approach for planning the pursuer trajectory. [^1] https://github.com/RajPShinde/Dynamic-Window-Approach
III-C Factor Graphs
This paper proposes a factor graph based solution to determine the optimal motion of pursuers. In factor graphs, nodes represent variables, such as a robot’s pose, while edges represent relationships between these variables, including kinematic constraints and measurements. An optimization cost is defined based on minimizing the costs associated with the edges of the factor graph. The optimization problem is formulated as a maximum a posteriori (MAP) inference problem and is implemented using GTSAM [29]. The Levenberg–Marquardt (LM) algorithm [30] is employed as the optimizer, providing fast convergence. The factor graph representation of the PE problem is adaptable and can be designed to enable the pursuers to move within the search space to actively pursue the evader.
@@ III-C Factor Graphs
Figure 1: The factor graph representation for the PE problem. Green nodes represent the variables associated with the poses of the pursuers. Red nodes are poses of evaders. Squares represent the constraints between the variables, including motion models (e.g. $f_{1}^{dp_{1}}$), measurements of the robots (e.g. $f_{0}^{mp_{1}}$) or landmarks (e.g. $f_{1}^{m_{1}o_{1}}$).
Fig. 1 illustrates the factor graph of the PE problem. Variables that need to be estimated or optimized are represented as nodes inside circles (green for pursuers and red for the evader). $x_{t}^{p_{i}}$ denotes the pose of pursuer $i$ at time $t$, while $x_{t}^{q}$ represents the pose of the evader at time $t$ and $T$ shows the current time step.
The overall cost function can be defined as
$J(x^{q}_{0:T},x^{p}_{0:T})$ $=\sum_{t=0}^{T}\Big{(}e_{t}^{qd}+\sum_{j=1}^{N_{P}}e_{t}^{mp_{j}}\quad+\sum_{j=1}^{N_{p}}\sum_{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{N_{P}}e_{t}^{cp_{j}}$
$+\sum_{j=1}^{N_{p}}\sum_{i=1}^{N_{o}}e_{t}^{o_{i}p_{j}}+e^{m_{j}o_{i}}+\sum_{j=1}^{N_{p}}e_{t}^{p_{j}d}+\sum_{j=1}^{N_{p}}e_{t}^{p_{j}mov}\Big{)}$ (4)
where, ||.||||.||| | . | | denotes the Mahalanobis distance [31], and $\Omega$ represents the information matrices of the variables. Here, $e^{qd}_{t}$ represents the cost associated with the dynamic factor of the evader, while $e_{t}^{p_{j}d}$ denotes the dynamic factor cost of the $j^{th}$ pursuer. The measurement factor costs are denoted by $e^{mp_{j}}_{t}$ and $e_{t}^{o_{i}p_{j}}$. Specifically, $e^{mp_{j}}_{t}$ indicates the measurement factor cost between the $j^{th}$ pursuer and the evader, whereas $e_{t}^{m_{j}o_{i}}$ represents the measurement factor cost between the $i^{th}$ obstacle and the $j^{th}$ pursuer. $e_{t}^{p_{j}mov}$ represents the movement cost for the $j^{th}$ pursuer, minimizing the distance between the evader’s estimated position and the $j^{th}$ pursuer. The terms $e_{t}^{o_{i}p_{j}}$ and $e_{t}^{cp_{j}}$ correspond to the cost of the obstacle avoidance and the collision avoidance factor, respectively. The problem consists of seven types of factors, described below:
III-C1 Dynamic Factor of Evader

### v2

@@ Abstract
keywords:
Multi-Robot , Factor Graph , Pursuit-Evasion , Uncertainty , Tracking
journal: Robots and Autonumous Systems
1 Introduction
@@
Several methods have been utilized to solve the PE problem. Classical methods use graph-based techniques, treating the PE problem as a search problem [6, 7]. Reinforcement learning (RL) approaches [8, 9] and game theory methods [10] have also been investigated. However, existing methods often lack an efficient representation of the relationships between pursuers, evaders, and their environment. For instance, if one pursuer loses its history of movements, retrieving that information can be challenging. Utilizing a graph-based approach can establish structural connections between entities, improving the overall understanding of the situation. In addition, the aforementioned methods do not account for uncertainties in predictions and do not facilitate message passing among robots. This limitation affects the scalability and versatility of PE variants.
In this paper, we present FG-PE, a method that employs factor graphs (FG) [11] to tackle the PE problem. Factor graphs provide a flexible framework capable of adapting to varying numbers of pursuers and obstacles. This approach addresses prediction uncertainty by guiding pursuers to minimize capture time, distance, and uncertainty, utilizing information from sensor measurements and past observations.
@@
The main contribution of the proposed method, FG-PE, lies in formulating a factor-graph-based solution to the pursuit–evasion problem. To the best of our knowledge, this is the first work to model the PE problem using a factor graph. FG-PE achieves up to a 100-fold reduction in capture time compared to benchmark methods, while requiring fewer measurements. This leads to lower memory usage and computational overhead, enabling FG-PE to outperform Constant Bearing (CB), Pure Pursuit (PP)  [12], and GUT [13]. The key features of FG-PE are as follows: (1) FG-PE can estimate the evader’s position and plan the pursuers while considering short time windows. This means that at each time step, it can determine the planned position for each pursuer for the subsequent time step. (2) FG-PE accounts for uncertainty in predictions using the probabilistic nature of factor graphs. It aims to minimize this uncertainty by providing a plan through factor graphs, which enable efficient representation and optimization in the presence of noise. Videos of the experiments are available on the website of the project: https://sites.google.com/view/pursuit-evasion. (3) FG-PE is easily extendable to multiple pursuers and allows for high obstacle-wise scalability. (4) FG-PE is robust against lost messages and performs well even when some messages are dropped within the graph structure.
2 Literature Review and Background
@@
The PE problem finds applications in various robotics scenarios, such as surveillance, search and rescue operations, and multiplayer games [14]. PE involves two groups of agents: pursuers and evaders. The evaders aim to reach a predefined goal, such as a specific location, object, or target area, while the pursuers try to prevent them from doing so. Key aspects of the PE problem include multi-agent tracking [15, 16] and path planning [2]. There are three main approaches solving the PE problem: search-based, learning-based, and graph-based approaches.
@@
A direct method for addressing the PE problem is to use search algorithms like Depth-First Search (DFS) [17] or Breadth-First Search (BFS) [18]. A critical aspect of the pursuit-evasion problem is the speed of the agents. Furhan et al. [19] introduced a method that accounts for uncertainties in the pursuers’ speed, which can lead to issues such as moving to unintended locations. A cooperative strategy has been developed for multiple pursuers to capture a high-speed evader in holonomic systems, addressing both formation and capture requirements [20]. Multi-pursuer pursuit–evasion games with a high-speed evader are analyzed using perfectly encircled formations and Apollonius circles, highlighting conditions for guaranteed capture as well as evader escape strategies [21].
@@
Several studies have derived optimal evasion strategies under specific pursuer behaviors. Makkapati et al. [29, 30] analyzed a two-pursuer, one-evader scenario and derived analytical strategies when the pursuers follow either constant-bearing or pure-pursuit trajectories. In a more general setting, Makkapati and Tsiotras [12] extended these results to multi-player pursuit–evasion problems by combining optimal evasion with task allocation among pursuers. These works provide important benchmarks for evaluating evader strategies, although they rely on the assumption of specific pursuer behaviors.
Several approaches have addressed the pursuit–evasion problem in the presence of external flow fields. One method approximates the flow using time-invariant affine functions [31]. Another line of work handles dynamic disturbances by defining reachable set inclusions, where level-set equations are solved to generate the pursuer’s reachable sets [32]. Reachable set–based formulations have also been extended to pursuit–evasion scenarios involving one evader and multiple pursuers [33]. Multi-pursuer, multi-evader differential games in three-dimensional environments with dynamic disturbances have been studied using partitioning strategies and reachability analysis, enabling the computation of time-optimal trajectories and strategies [34]. More recently, single-pursuer, single-evader games in stochastic flow fields have been analyzed using forward reachability and chance-constrained linear–quadratic formulations, solved via Gauss–Seidel iterations [35].
@@
There are several real-world applications that require the collaboration of multiple robots such as object transportation [36] and task allocation [37]. One approach to coordinating the necessary communication among robots is by utilizing graph representation. Factor graphs [11], described in Sec. 3.3 in detail, can represent complex relationships between agents concisely and are suitable for addressing a wide range of problems such as kinodynamic motion planning that fully considers whole-body dynamics and contacts [38]. Factor graphs have been widely used to solve various multi-robot problems, such as localization [39], path planning [40], and control [4].
2.1 Factors in Factor Graphs
Factor graphs are bipartite graphs that represent how a global function can be decomposed into a product of local functions, called factors [41]. Normally, given a set of variables $X=\{x_{1},x_{2},\dots,x_{n}\}$, a factor graph expresses the joint distribution as
$f(X)=\prod_{a\in\mathcal{F}}f_{a}(X_{a}),$ (1)
where $\mathcal{F}$ denotes the set of factors in the factor graph, and each factor $f_{a}$ is a local function depending only on a subset of variables $X_{a}\subseteq X$. Factor nodes are connected to variable nodes when the corresponding variable appears in the factor. Several types of factors appear frequently in robotics and estimation problems:
2.1.1 Unary factors
Unary factors depend on a single variable. They typically encode prior information or measurements about an individual variable:
$f(x_{i})\propto\exp\Bigg(-\frac{1}{2}\|h(x_{i})-z\|^{2}_{\Omega}\Bigg),i=1,2,...,n$ (2)
where $z$ is the measurement, $h(\cdot)$ is the measurement model. Here $||.||$ denotes the Mahalanobis distance [42], and $\Omega$ represents the information matrices of the variables. For the Gaussian distribution, $\Omega=\Sigma^{-1}$, $\Sigma$ is the covariance of the variable. A prior factor is a special case of a unary factor. It encodes prior knowledge about a variable and effectively anchors it to a known or assumed value while allowing for uncertainty. Mathematically, a prior factor constrains a variable $x$ to a prior mean $\mu$ with information matrix $\omega$ (covariance $\Sigma=\Omega^{-1}$), and is typically modeled as a Gaussian:
$f_{\text{prior}}(x)\propto\exp\Big(-\frac{1}{2}\|x-\mu\|_{\Omega}^{2}\Big),$ (3)
where
$\|x-\mu\|^{2}=(x-\mu)^{\top}\Omega(x-\mu)$ (4)
is the Mahalanobis distance between the variable $x$ and the prior mean $\mu$, weighted by the information matrices $\Omega$.
2.2 Pairwise factors
Pairwise factors involve two variables and are often used to express relative constraints between them:
$f(x_{i},x_{j})\propto\exp\Bigg(-\frac{1}{2}\|h(x_{i},x_{j})-z\|^{2}_{\Omega}\Bigg).$ (5)
These appear in odometry, landmark observations, or inter-robot measurements.
2.2.1 Higher-order factors
Higher-order factors involve more than two variables. They are used when the constraint naturally depends on a set of variables simultaneously, such as synchronization:
$f(x_{i},x_{j},x_{k})\propto\exp\Bigg(-\frac{1}{2}\|h(x_{i},x_{j},x_{k})-z\|^{2}_{\Omega}\Bigg).$ (6)
In this paper, we utilize factor graphs to address the PE problem. To the best of the authors’ knowledge, there is no existing work, that formulates the PE problem using a factor graph representation.
3 Proposed Method
In this section, the problem statement and the proposed solution are presented. Table 1 lists the notations used; $q$ is used for the evader, and $p_{j}$ is used for the $j^{\text{th}}$ pursuer. Factors are denoted by $f$, with superscripts indicating the type of constraint, such as dynamics (i.e. $f^{d}$), measurement (i.e. $f^{m}$), planning (i.e. $f^{mov}$), collision avoidance (i.e. $f^{c}$), and obstacle avoidance (i.e. $f^{o}$). Each factor $f$ corresponds to an error term $e$, which represents the cost minimized during optimization. The same notation is used for the cost. For factors and costs, after the constraint type, further symbols indicate if the factor/cost is associated with a pursuer, an evader, or an obstacle. The indices $i$ and $j$ denote the identifiers of the corresponding pursuers or obstacles, while the subscript $t$ represents the discrete time step. Note, there is only one evader.
Symbol Description
Variables
@@ 3 Proposed Method
$f^{mp_{j}}_{t}$ Measurement factor between $j^{th}$ pursuer and the evader at time $t$
$f^{mo_{ji}}_{t}$ Measurement factor between $j^{th}$ pursuer and $i^{th}$ obstacle at time $t$
$f^{movp_{j}}_{t}$ Planning factor of the $j^{th}$ pursuer at time $t$
$f^{cp_{ji}}_{t}$ Collision avoidance factor between pursuer $j$ and pursuer $i$ at time $t$
$f^{op_{ij}}_{t}$ Obstacle avoidance factor between $j^{th}$ pursuer and $i^{th}$ obstacle at time $t$
Costs (Error Terms)
$e^{dq}_{t}$ Cost of dynamic factor of the evader
$e^{dp_{j}}_{t}$ Cost of dynamic factor of the $j^{th}$ pursuer
$e^{mp_{j}}_{t}$ Cost of measurement factor between $j^{th}$ pursuer and evader
$e^{mo_{ji}}_{t}$ Cost of measurement factor between $j^{th}$ pursuer and $i^{th}$ obstacle
$e^{movp_{j}}_{t}$ Cost of planning factor for pursuer $j$
$e^{cp_{ji}}_{t}$ Cost of collision avoidance between pursuer $j$ and pursuer $i$
$e^{op_{ji}}_{t}$ Cost of obstacle avoidance factor between pursuer $j$ and obstacle $i$
Table 1: Summary of notation. Variables represent unknown quantities to be estimated, factors encode probabilistic or planning constraints, and costs correspond to the residual errors associated with each factor.
3.1 Problem Statement
@@
[…]
@@
3.2 Path Planning of Evader
The evader aims to reach a predefined goal, known only to itself. The evader employs the Dynamic Window Approach (DWA), a popular method in robotics for motion planning and obstacle avoidance [43]. DWA is effective in scenarios where a robot navigates through moving obstacles. It dynamically adjusts the robot’s velocity and steering commands based on real-time sensory input and desired goal states. By continuously evaluating and updating potential future trajectories within a dynamically defined window, the evader can efficiently navigate toward its goal while avoiding obstacles and pursuers. The evader iterates through all possible trajectories based on angular velocity and linear velocity, selecting the best movement for the next time step. In this paper, we utilize a modification of the DWA algorithm [^1] https://github.com/RajPShinde/Dynamic-Window-Approach for planning the evader’s trajectory.
3.3 Factor Graphs
This paper proposes a factor graph based solution to determine the optimal motion of pursuers. In factor graphs, nodes represent variables, such as a robot’s pose, while edges represent relationships between these variables, including kinematic constraints and measurements.
An optimization cost is defined based on minimizing the costs associated with the edges of the factor graph. The optimization problem is formulated as a maximum a posteriori (MAP) [11] inference problem and is implemented using the Georgia Tech Smoothing and Mapping (GTSAM) library [44]. It represents inference problems using factor graphs and Bayesian networks, which provide a structured framework for optimization. These problems are then solved efficiently using sparse matrix factorization techniques to compute an optimal solution. The Levenberg–Marquardt (LM) algorithm [45] is employed as the optimizer, providing fast convergence. LM is a widely used iterative method for solving nonlinear least-squares problems, combining the advantages of the Gauss–Newton algorithm and gradient descent. The factor graph representation of the PE problem is shown in Fig. 1 and it is adaptable and can be designed to enable the pursuers to move within the search space to actively pursue the evader.
@@ 3.3 Factor Graphs
Figure 1: The factor graph representation for the PE problem. Green nodes represent the variables associated with the poses of the pursuers. Red nodes are poses of evaders. Squares represent the constraints between the variables, including motion models (e.g. $f_{1}^{dp_{1}}$), measurements of the robots (e.g. $f_{0}^{mp_{1}}$) or landmarks (e.g. $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{1}^{mo_{11}}$}).
The overall cost function at time $T$ can be defined as
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{J}}(x^{q}_{0:T},x^{p}_{0:T})$ $=\sum_{t=0}^{T}\Big(e_{t}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}dq}}+\sum_{j=1}^{N_{P}}e_{t}^{mp_{j}}\quad+\sum_{j=1}^{N_{p}}\sum_{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{N_{P}}e_{t}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}cp_{ji}}}$
$+\sum_{j=1}^{N_{p}}\sum_{i=1}^{N_{o}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(}e_{t}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}op_{ji}}}+e^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}mo_{ji}}}_{t}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0})}+\sum_{j=1}^{N_{p}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}e_{t}^{dp_{j}}}+\sum_{j=1}^{N_{p}}e_{t}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}movp_{j}}}\Big).$ (10)
Here, $e^{dq}_{t}$ represents the cost associated with the dynamic factor of the evader, while $e_{t}^{dp_{j}}$ denotes the dynamic factor cost of the $j^{th}$ pursuer. The measurement factor costs are denoted by $e^{mp_{j}}_{t}$ and $e_{t}^{op_{ij}}$. Specifically, $e^{mp_{j}}_{t}$ indicates the measurement factor cost between the $j^{th}$ pursuer and the evader, whereas $e_{t}^{mo_{ji}}$ represents the measurement factor cost between the $i^{th}$ obstacle and the $j^{th}$ pursuer. $e_{t}^{movp_{j}}$ represents the movement cost for the $j^{th}$ pursuer, minimizing the distance between the evader’s estimated position and the $j^{th}$ pursuer. The terms $e_{t}^{op_{ji}}$ and $e_{t}^{cp_{ji}}$ correspond to the cost of the obstacle avoidance and the collision avoidance factor, respectively. The overall cost function is iterating through all factors and the aim is to minimize the overall cost function over all factors in the formulated factor graph at time $T$. The problem consists of seven types of factors, described below:
3.3.1 Dynamic Factor of Evader